# CURE: A dataset for Clinical Understanding & Retrieval Evaluation

### v3

@@ 3 Dataset Statistics
Figure 1: Query Type Distribution across the domains in CURE
@@
Table 5: Descriptive statics for all domains in CURE. # Avg. Q Len: average number of tokens per query for English (en), Spanish (es), & French (fr); # Avg. P Len: average number of tokens per passage. Tokens are delimited by whitespace.
4 Retrieval Baselines
To demonstrate the effectiveness of our test dataset for retrieval we present results from three different baselines. BM25 serves as our representative for a sparse retrieval baseline. For a dense retrieval baseline we choose a GTE Multilingual Base Zhang et al. (2024) as our representative for models with fewer than 500M parameters and NV Embed V2 as our representative for models with greater than 1B parameters Lee et al. (2024). We will make these baselines available for reproducibility.
Results are presented in Table 6. We used the MTEB evaluation framework Muennighoff et al. (2022) to compute the results. As expected, the the largest dense performance, that of NV-Embed, bests that of the smaller parameter dense model, GTE Multilingual Base, which in turn outstrips the sparse model. Across all retrievers, performance on the monolingual conditions is better than model across cross-lingual conditions of CURE.
@@ 4 Retrieval Baselines
BM-25 Lù (2024) 0.355 0.018 0.012 0.523 0.044 0.084
mGTE Base (304M) Zhang et al. (2024) 0.558 0.528 0.481 0.714 0.623 0.643
NV Embed v2 (7B) Lee et al. (2024) 0.651 0.603 0.604 0.786 0.738 0.740
Table 6: Sparse and dense baseline results on CURE.
5 Related Test Datasets
In this section, we review widely used biomedical test datasets considering their utility in evaluating ad hoc retrieval systems intended for use by healthcare professionals.
Unlike CURE, many commonly used test datasets that can be leveraged to evaluate retrieval do not reflect the broad range of concerns of medical practitioners. Datasets such as TREC-COVID Roberts et al. (2021) and SciFacts Wadden et al. (2020) have a narrow focus with corpora drawn from CORD-19, a collection of scientific articles on COVID-19 Roberts et al. (2021). The Nutrition Facts Corpus dataset Boteva et al. (2016) is similar with the corpus is constructed from academic texts in the field of nutrition. Other datasets are too broad, extending beyond far the field of medicine. Scifacts Wadden et al. (2020), for example, has a corpus that includes articles drawn from a broad range of scientific disciplines. BioASQ Krithara et al. (2023) as used in BEIR Thakur et al. (2021) has a domain focus most similar to CURE with a corpus constructed from PubMed/MedLine articles. However, queries in the dataset are more representative of the information needs of biomedical researchers than of healthcare providers Krithara et al. (2023).
@@ 5 Related Test Datasets
Nutrition Facts Boteva et al. (2016) Titles, Descriptions etc. [1] Nutrition 3.8 3237 12334 PubMed Abstracts
TREC-COVID Roberts et al. (2021) Natural questions COVID 50 493.5 24673 CORD-19 Papers
SciDocs Cohan et al. (2020) Article titles [1] Science 1000 29.9 29928
SciFacts Wadden et al. (2020) Assertions COVID 1119 1.1 1258 CORD-19 Papers
BioASQ Krithara et al. (2023) Natural questions Biomedicine 5389 9.7 52413 PubMed
@@
Table 7: Comparison of related retrieval collections. Human Labels indicates human-generated labels (vs. synthetically generated labels); CLIR indicates whether the collection supports cross-lingual retrieval; Avg # Labels/Q is the average number of labels provided per query; Total # Labels is the total number of human labels (both positive and negative) across all splits. 1Nutrition Facts & SciDocs construct queries from article titles; 2We translate human-generated queries; 3We also annotate query-passage pairs using LLMs as described in subsection 2.3.
@@
Furthermore, many biomedical retrieval datasets are repurposed from datasets originally intended to evaluate systems on downstream tasks other than ranking. As a consequence, queries may not be in forms representative of those posed by actual users. For example, in the Nutrition Facts Corpus dataset Boteva et al. (2016), queries are constructed from article titles and descriptions. The SciDocs dataset Wadden et al. (2020), originally designed to test citation prediction and repurposed for retrieval evaluation in BEIR Wadden et al. (2020), uses paper titles as queries. Similarly, the Scifacts dataset as as used in BEIR Wadden et al. (2020), originally designed to evaluate fact-checking, uses assertions rather than interrogatives as queries. In contrast CURE includes natural language queries formulated by health practitioners.
@@
The relevance judgments in some test retrieval datasets are also limited. For example, in both SciDocs Wadden et al. (2020) and Nutrition Facts Corpus Boteva et al. (2016) relevance is inferred from citations rather than explicitly judged. In CURE, we include explicit human relevance judgments and mitigate bias introduced by relatively shallow candidate pools by using an LLM to fill in missing relevance judgments.
6 Limitations and Future Work
@@
CURE currently covers a limited number of medical domains. As performance in one domain may not be indicative of performance in another domain, in future iterations we would like to expand CURE’s coverage. We would also like to further diversify the type of queries represented in CURE, with a greater focus on queries that require reasoning, particularly within the context of clinical decision making.
@@
Although CURE supports the evaluation of search in academic literature, in a point-of-care setting, a medical practitioner may consult a range of different types of documents, including educational material or Electronic Health Records. In the future, we would like to develop test datasets that enable evaluation of retrieval systems across different types of documents.
7 Conclusion
@@
In this paper we present CURE, a test dataset to evaluate retrieval systems in healthcare settings. CURE was built with a team of healthcare professionals including nurses and doctors, who generated questions and identified relevant passages from biomedical literature. Its corpus was augmented by mining an index of 7777M passages extracted from biomedical articles. To mitigate bias, large candidate pools were constructed and an LLM was used to annotate passages for relevance. The final test dataset consists of 2000 queries evenly distributed over 10 medical domains, and is available in a monolingual (English) and two cross-lingual conditions (French/Spanish $\rightarrow$ English). We also present the performance of three retrievers on this dataset as baselines.
@@
There are few if any test datasets designed to evaluate retrieval systems intended for use by medical practitioners in a point-of-care setting. At Clinia, we use CURE to benchmark our existing retrieval solutions and guide their development. In making CURE freely available for evaluation, we aim to foster research and the development of reliable retrieval tools for healthcare practitioners.
Acknowledgments
@@
We want to highlight our team of annotators, particularly Dr. Arturo Vela Lasagabaster, MSc, MD; Dr. Leila Cattelan, MD; Dr. Natalia Cullell, MD, PhD; and Mr. Noah Tremblay Taillon for their meticulous work. We also thank our CTO, Etienne Soulard-Geoffrion, and the rest of Clinia’s executive team for their belief in responsible ML development and their support for this initiative. Finally, we thank the Governments of Canada and Quebec for their financial support in advancing technological innovation.

### v4

@@ 3. Dataset Statistics
Figure 1. Query Type Distribution across the domains in CURE
@@
Table 6. Descriptive statistics for all domains in CURE. # Avg. Q Len: average number of tokens per query for English (en), Spanish (es), & French (fr); # Avg. P Len: average number of tokens per passage. Tokens are delimited by whitespace.
4. Retrieval Baselines
To demonstrate the effectiveness of our test dataset for retrieval, we present results from four different baselines. BM25 serves as our representative sparse retrieval baseline. For a dense retrieval baseline, we choose MedEmbed Base (Balachandran, 2024) as a representative of a domain-specific model, GTE Multilingual Base (Zhang et al., 2024) as our representative for models with fewer than 500M parameters, and NV Embed V2 as our representative for models with more than 1B parameters (Lee et al., 2024b).
The results are presented in Table 7. We used the MTEB evaluation framework (Muennighoff et al., 2022) to compute the results. As expected, the largest dense model, NV-Embed, outperforms the smaller dense models, GTE Multilingual Base and MedEmbed Base, which in turn outperforms the sparse BM25 model. Across all retrievers, the performance under monolingual conditions is better than the model under cross-lingual conditions of CURE.
@@ 4. Retrieval Baselines
BM-25 (Lù, 2024) 0.355 - - 0.523 - -
MedEmbed-base-v0.1 (Balachandran, 2024) 0.515 0.193 0.144 0.690 0.368 0.308
mGTE Base (304M) (Zhang et al., 2024) 0.558 0.528 0.481 0.714 0.623 0.643
NV Embed v2 (7B) (Lee et al., 2024b) 0.651 0.603 0.604 0.786 0.738 0.740
Table 7. Sparse and dense baseline results on CURE.
There are many compelling retrieval models available online, making it challenging to select just a few for presentation. As an additional reference, results for a subset of well-known multilingual dense embedders, both open source and proprietary, are provided in Table 8.
For readers interested in evaluating a broader range of models, a comprehensive and up-to-date overview of additional results is available on the MTEB Leaderboard (Muennighoff et al., 2022). Results specific to CURE can be viewed by selecting the MTEB(Medical, v1) benchmark and filtering by task.
The MTEB(Medical, v1) benchmark, which not only includes CURE but also features tasks from medical settings distinct from point-of-care, was proposed by Clinia to facilitate the evaluation of new models. As such, Clinia remains committed to maintaining it by continuously evaluating and uploading results for new models as they become available.
NDCG@10 Recall@100
Model Name en-en fr-en es-en en-en fr-en es-en
paraphrase-multilingual-MiniLM-L12-v2 (Reimers and Gurevych, 2019) 0.309 0.203 0.225 0.481 0.337 0.366
multilingual-e5-small (Wang et al., 2024) 0.504 0.202 0.163 0.648 0.332 0.281
multilingual-e5-base (Wang et al., 2024) 0.553 0.257 0.194 0.697 0.407 0.340
snowflake-arctic-embed-m-v2.0 (Yu et al., 2024) 0.559 0.485 0.491 0.715 0.632 0.646
snowflake-arctic-embed-l-v2.0 (Yu et al., 2024) 0.573 0.518 0.522 0.714 0.650 0.657
jina-embeddings-v3 (Sturua et al., 2024) 0.484 0.428 0.441 0.656 0.581 0.607
google-text-multilingual-embedding-002 (Lee et al., 2024a) 0.510 0.395 0.403 0.689 0.576 0.591
cohere-embed-multilingual-v3.0 (Cohere, 2023) 0.587 0.522 0.527 0.730 0.576 0.663
openai-text-embedding-3-small (OpenAI, 2024b) 0.520 0.430 0.427 0.697 0.611 0.614
openai-text-embedding-3-large (OpenAI, 2024b) 0.598 0.559 0.554 0.751 0.712 0.713
Table 8. Additional results for various multilingual dense embedders on CURE.
5. Related Test Datasets
In this section, we review widely used biomedical test datasets considering their utility in evaluating ad hoc retrieval systems intended for healthcare professionals.
Unlike CURE, many commonly used test datasets for evaluating retrieval do not reflect the wide range of concerns of medical professionals. Datasets such as TREC-COVID (Roberts et al., 2021) and SciFacts (Wadden et al., 2020) have a narrow focus, with corpora drawn from CORD-19, a collection of scientific articles on COVID-19 (Roberts et al., 2021). The Nutrition Facts Corpus dataset (Boteva et al., 2016) is similar, as its corpus is constructed from academic texts in the field of nutrition. Other datasets are too broad, extending beyond the field of medicine. Scifacts (Wadden et al., 2020), for example, has a corpus that includes articles drawn from a wide range of scientific disciplines. BioASQ (Krithara et al., 2023), as used in BEIR, (Thakur et al., 2021) has a domain focus more similar to CURE, with a corpus constructed from PubMed/MedLine articles. However, queries in BioASQ are broadly representative of the information needs of biomedical scientists, rather than exclusively those of healthcare providers (Krithara et al., 2023). In contrast, the queries in CURE were generated by medical professionals and explicitly validated as being typical of those asked by medical providers in a point of care setting. Of the queries generated, 97.63% of the queries generated for CURE were identified by 2 or more out of 3 annotators as being aligned with the types of questions likely to be asked by a health practitioner seeking research articles in a point-of-care setting.
@@ 5. Related Test Datasets
Nutrition Facts (Boteva et al., 2016) Titles, Descriptions etc. [1] Nutrition 3.8 3237 12334 PubMed Abstracts
TREC-COVID (Roberts et al., 2021) Natural questions COVID 50 493.5 24673 CORD-19 Papers
SciDocs (Cohan et al., 2020) Article titles [1] Science 1000 29.9 29928
SciFacts (Wadden et al., 2020) Assertions COVID 1119 1.1 1258 CORD-19 Papers
BioASQ (Krithara et al., 2023) Natural questions Biomedicine 5389 9.7 52413 PubMed
@@
Table 9. Comparison of related retrieval collections. Human Labels indicates human-generated labels (vs. synthetically generated labels); CLIR indicates whether the collection supports cross-lingual retrieval; Avg # Labels/Q is the average number of labels provided per query; Total # Labels is the total number of human labels (both positive and negative) across all splits. 1Nutrition Facts & SciDocs construct queries from article titles; 2We translate human-generated queries; 3We also annotate query-passage pairs using LLMs as described in subsection 2.3.
@@
Furthermore, many biomedical retrieval datasets are repurposed from datasets originally intended to evaluate systems on downstream tasks other than retrieval. As a consequence, queries may not be in forms representative of those posed by actual users. For example, in the Nutrition Facts Corpus dataset (Boteva et al., 2016), queries are constructed from article titles and descriptions. The SciDocs dataset (Wadden et al., 2020), originally designed to test citation prediction and repurposed for retrieval evaluation in BEIR (Wadden et al., 2020), uses titles of papers as queries. Similarly, the Scifacts dataset, as used in BEIR (Wadden et al., 2020), originally designed to evaluate fact-checking, uses assertions rather than interrogatives as queries. In contrast, CURE includes queries in natural language formulated by healthcare professionals.
@@
Relevance judgments in some test retrieval datasets are also limited. For example, in both SciDocs (Wadden et al., 2020) and Nutrition Facts Corpus (Boteva et al., 2016) relevance is inferred from citations rather than explicitly judged. In CURE, we include explicit human relevance judgments and mitigate the bias introduced by relatively shallow candidate pools by using an LLM to fill in missing relevance judgments.
6. Limitations and Future Work
@@
CURE currently covers a limited number of medical domains. As performance in one domain may not be indicative of performance in another, in future iterations we would like to expand CURE’s coverage. We would also like to further diversify the type of queries represented in CURE, with a greater focus on queries that require reasoning, particularly within the context of clinical decision-making.
@@
Although CURE supports the evaluation of search in academic literature, in a point-of-care setting a medical practitioner may consult a range of different types of documents, including educational materials or Electronic Health Records. In the future, we would like to develop test datasets that enable the evaluation of retrieval systems across different types of documents.
7. Conclusion
@@
In this paper, we present CURE, a test dataset to evaluate retrieval systems in healthcare settings. CURE was built with a team of healthcare professionals, including nurses and doctors, who generated questions and identified relevant passages from biomedical literature. Its corpus was augmented by mining an index of 7777M passages extracted from biomedical articles. To mitigate bias, large candidate pools were constructed and an LLM was used to annotate passages for relevance. The final test dataset consists of 2,000 queries evenly distributed over 10 medical domains and is available in a monolingual (English) and two cross-lingual conditions (French/Spanish to English). We also present the performance of different retrievers on this dataset as baselines.
@@
There are few, if any, test datasets designed to evaluate retrieval systems intended for use by medical practitioners in a point-of-care setting. At Clinia, we use CURE to benchmark our existing retrieval solutions and guide their development. In making CURE freely available for evaluation, we aim to foster research and the development of reliable retrieval tools for healthcare practitioners.
Acknowledgements.
@@
We want to highlight our extensive team of annotators, particularly Dr. Arturo Vela Lasagabaster, MSc, MD; Dr. Leila Cattelan, MD; Dr. Natalia Cullell, MD, PhD; and Mr. Noah Tremblay Taillon for their meticulous work. We also extend our gratitude to our CTO, Etienne Soulard-Geoffrion, and the rest of Clinia’s executive team for their belief in responsible ML development and their support for this initiative.
Appendix A Annotation Interface
A screenshot of the interface used to collect annotations is presented in Figure 2
Figure 2. Screenshot of Argilla Interface used for Annotation Generation
Appendix B Prompt used for Relevance Judgments
You are a { clinical_role }, who consults academic journals to ensure that you provide care that is in accordance with best practices. Here, you will be presented a query you constructed along with a retrieved passage from a scientific article. Your task is to assign a score on an integer scale of 0 to 2 based on the degree to which the passage contains an actionable answer to your query:
2: Assign a score of 2 if the passage is dedicated to the query and contains the exact answer
1: Assign a score of 1 if the passage has some answer for the query but the answer is a bit unclear, or hidden amongst extraneous information
0: Assign a score of 0 if the passage seems related to the query but does not answer it OR if the passage has nothing to do with the query
Note: The article title is provided to help you contextualize the passage.
# Important Instruction
Split this task into steps:
1. Identify the underlying information need of the search from the perspective of a trained { profession } seeking to provide informed patient care.
2. Measure how well the passage satisfies the information need expressed by the query. Carefully consider details in both the query and the passage.
3. Before scoring the passage, write a careful explanation justifying your score that includes a detailed rationale.
4. Provide each score in the format of a JSON with the following structure:
"explanation": "Careful explanation justifying your score that includes a detailed rationale",
"score": "Your final assigned score based on the explanation. One of: 0, 1, 2"
Query: { query }
Passage: { passage }
Article Title: { article_title }
# Examples
{query_examples}